Acton Company has two products: A and B. The annual production and sales of Product A is 800 units and of Product B is 500 units. The company has traditionally used direct labor-hours as the basis for applying all manufacturing overhead to products. The total estimated overhead for next period is $92,023. The company is considering switching to an activity-based costing system for the purpose of computing unit product costs for external reports. The new activity-based costing system would have three overhead activity cost pools—Activity 1, Activity 2, and General Factory—with estimated overhead costs and expected activity as follows:

Estimated

Activity

Activity

Overhead

Expected Activity

Cost Pool

Measure

Costs

Product A

Product B

Total

Activity 1

number of setups

$14,487

500

600

1,100

Activity 2

machine hours

$64,800

2,500

500

3,000

General Factory

direct labor hours

$12,736

240

100

340

   Total

$92,023

The predetermined overhead rate under the traditional costing system is closest to:

$37.46

$21.60

$13.17

$270.66

The overhead cost per unit of Product B under the traditional costing system

$54.13

$7.49

$4.32

$2.63

.   The predetermined overhead rate (i.e., activity rate) for Activity 1 under the activity-based costing system is closest to:

$28.97

$13.17

$83.66

$24.15

The overhead cost per unit of Product A under the activity-based costing system is closest to:

$86.97

$70.79

$81.20

$11.24
Answer #1

1) Predetermine overhead rate = 92023/340 = 270.66

So answer is d) $270.66

2) Product B Overhead cost per unit = (270.66*100)/500 = 54.13

So answer is a) $54.13

3) Overhead rate for activity 1 = 14487/1100 = 13.17

So answer is b) $13.17

4) Product A overhead cost per unit

Product A
Activity A 13.17*500 = 6585
Activity B 21.60*2500 = 54000
General factory 37.46*240 = 8990.40
Total overhead 69575.40
Unit 800
Overhead cost per unit 86.97

So answer is a) $86.97